# Déjà Vu: Efficient Video-Language Query Engine with Learning-based Inter-Frame Computation Reuse

### v1

@@ Abstract.
Recently, Video-Language Models (VideoLMs) have demonstrated remarkable capabilities, offering significant potential for flexible and powerful video query systems. These models typically rely on Vision Transformers (ViTs), which process video frames individually to extract visual embeddings. However, generating embeddings for large-scale videos requires ViT inferencing across numerous frames, posing a major hurdle to real-world deployment and necessitating solutions for integration into scalable video data management systems. This paper introduces Déjà Vu, a video-language query engine that accelerates ViT-based VideoLMs by reusing computations across consecutive frames. At its core is ReuseViT, a modified ViT model specifically designed for VideoLM tasks, which learns to detect inter-frame reuse opportunities, striking an effective balance between accuracy and reuse. Although ReuseViT significantly reduces computation, these savings do not directly translate into performance gains on GPUs. To overcome this, Déjà Vu integrates memory-compute joint compaction techniques that convert the FLOP savings into tangible performance gains. Evaluations on three VideoLM tasks show that Déjà Vu accelerates embedding generation by up to a 2.64$\times$ within a 2% error bound, dramatically enhancing the practicality of VideoLMs for large-scale video analytics.
@@
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
The source code, data, and/or other artifacts have been made available at https://github.com/casys-kaist/DejaVu.
@@ 1. Introduction
Video data is pervasive in today’s world, playing a critical role in various applications such as autonomous machines (Deb et al., 2018; Gawande et al., 2020; Hadidi et al., 2018), education (Gan et al., 2023), healthcare (Sanderson and Matuszewski, 2022), security and surveillance (Li et al., 2019; Toor et al., 2019), e-commerce (Zhang et al., 2020a, 2022), and many others (Mun et al., 2017; Li et al., 2022d). Recently, Video-Language Models (VideoLMs) have demonstrated remarkable capabilities, offering significant potential for flexible and powerful video analytics tasks (Colas et al., 2019; Gupta and Demner-Fushman, 2022; Zhao et al., 2023; Ju et al., 2022; Yang et al., 2021, 2022a). While these powerful algorithmic advances promise enormous potential, the massive computational demand poses a significant hurdle for the widespread adoption of VideoLMs for video query systems.
@@
Figure 1 provides an overview of video-language query systems, where ViTs process video frames to extract visual embeddings, which are then fed into query-specific operations such as video retrieval, video question answering, and video question grounding. The critical role of ViTs in this process emphasizes the importance of improving ViT inference efficiency to unleash the potential of VideoLMs on building video-language query systems.
@@
ReuseViT: Reuse-enabling ViT model. We develop a modified ViT model, referred to as ReuseViT, that reuses precomputed values for similar tokens between frames. To maximize reuse opportunities, we devise a frame referencing strategy that determines inter-frame reference chains for computation reuse. ReuseViT contains Decision Layers that dynamically determine when to reuse computations by processing multiple inputs such as cosine similarity, attention scores, codec metadata, and reference types, allowing it to learn the importance and interactions of these hints automatically. Additionally, it employs Restoration Layers that learn to calibrate the reused computation values based on the changes occurred in the current frame.
@@
To demonstrate the effectiveness of Déjà Vu, we perform evaluations using three different VideoLMs – (1) Video Retrieval, (2) Video Question Answering, and (3) Video Question Grounding – on their respective datasets. We observe that Déjà Vu achieves throughput improvement of 1.81$\times$, 2.64$\times$, and 2.54$\times$ for the three tasks, respectively, within 2% error bound. Under the same accuracy constraints, the state-of-the-art systems using inter-frame computation reuse offer only 1.32$\times$, 2.08$\times$, 2.20$\times$ speedups for the respective tasks.
@@ 3.1. Frame Selection for Computation Reuse
[…]
@@ 3.2. Layer Selection for Reuse
Figure 5. FLOPs breakdown of core computations within a single encoder layer of vision transformers at different scale.
@@
Computation patterns of layers. Besides computational cost, we consider computation patterns of each layer. The QKV projection and FFN operations are applied independently to each token, without inter-token dependencies, suiting them for computation reuse. In contrast, the self-attention operation involves interactions among all tokens, making reuse more challenging and potentially affecting model output. Given these considerations, we focus on reusing computations in the QKV projection and FFN layers, targeting the most computationally intensive and token-independent operations.
@@ 3.3. Criteria for Computations Reuse
(4) $\mathcal{M}_{i}$ =$=\begin{cases}1,&\text{if }d_{i}>0\\ 0,&\text{otherwise}\end{cases},\quad i=1,\ldots,N$
@@
(5) $C$ $={T^{\text{cur}}_{i}\mid\mathcal{M}_{i}=0},\quad\text{(re\text@underline{c}ompute tokens)}$)
(6) $R$ $={T^{\text{cur}}_{i}\mid\mathcal{M}_{i}=1},\quad\text{(\text@underline{r}euse tokens)}$)
@@ 6.3. Online Inference
Once deployed, Déjà Vu employs layerwise scheduling and memory compaction to translate the FLOP savings from ReuseViT into practical GPU throughput. While ReuseViT reduces the token-level computational load, effective scheduling and compaction routines ensure these savings materialize in actual GPU execution time which is a key requirement for large-scale query systems.
@@
By blending standard ViT-based embedding generation with lightweight, learned modules for inter-frame reuse, Déjà Vu delivers on-demand embeddings for a variety of queries. Its approach avoids repeated backbone retraining, recycles partial results across neighboring frames, and capitalizes on compaction techniques to maximize GPU efficiency. These design choices not only make Déjà Vu robust to queries demanding high accuracy but also ensure practical deployment for both offline and streaming scenarios. As a result, Déjà Vu opens the door to scalable, open-ended video-language applications without the prohibitive compute costs usually associated with large-scale video analytics.
@@ 7.1. Methodology
End tasks. To evaluate the effectiveness of Déjà Vu, we select three distinct Video-Language Models (VideoLMs), each targetting a different task. All these models use embeddings extracted from CLIP (Radford et al., 2021) (ViT) as their visual backbone. We focus on:
  • Video retrieval. We use CLIP4Clip (Luo et al., 2022) for video retrieval and evaluate on MSR-VTT (Xu et al., 2016) dataset, which consists of 10,000 video clips annotated with textual descriptions, typically 10 to 30 seconds long. We report top-5 recall as the accuracy metric.
  • Video question answering. We use FrozenBiLM (Yang et al., 2022a) for video question answering (or video QA), a task where the goal is to answer queries based on video content. We evaluate on How2QA (Li et al., 2020a), which standardizes input to 60 second clips drawn from longer instructional videos with 44,007 question-answer pairs, and report multiple-choice accuracy as the metric.
  • Video Question Grounding. Video question grounding (or video grounded QA) requires the model to predict the relevant video segment along its answer. We use TempCLIP (Xiao et al., 2024) and evaluate it on NExT-GQA (Xiao et al., 2024), with longer, untrimmed videos averaging around 45 seconds, with short answer segments requiring fine-grained temporal grounding. We report GQA@Accuracy, the percentage of correct answers where the predicted video segment overlaps the ground-truth segment.
To ensure a fair comparison, all baselines sample frames at 2 frames per second, which aligns with common practice in recent video-language models (Piergiovanni et al., 2023; Arnab et al., 2021; Bulat et al., 2021; Zhang et al., 2023; Wang et al., 2024a; Luo et al., 2022; Yang et al., 2022a; Xiao et al., 2024). Throughout all tasks, we use the standard train, val, and test splits provided by each dataset and adhere to the recommended evaluation protocols.
Baselines. As no prior methods directly accelerate VideoLMs, our baselines include temporal ViT acceleration approaches (conceptually closest to Déjà Vu’s strategy) and a representative image-based method for contrast. We evaluate three acceleration methods:
  • CMC. CMC (Song et al., 2024) uses a fixed threshold value on the mean squared error (MSE) to identify unchanged tokens suitable for reuse across frames. It leverages similarity computations provided by specialized video codec hardware and was originally evaluated on action recognition tasks.
  • Eventful Transformer. Eventful Transformer (Dutson et al., 2023) applies a static, layer-wise policy that fixes the number of tokens to recompute, regardless of changes in video content. While simple and efficient, this rigid strategy might not adapt effectively to varying temporal content across frames, potentially causing error accumulation in dynamic scenes.
  • DiffRate. DiffRate (Chen et al., 2023) is a representative baseline for images, combining token pruning and merging within an image. It automatically determines the tradeoff between accuracy and efficiency through backpropagation. Though originally evaluated on image classification, we adapted the policy for VLP models.
@@ 7.3. Tradeoff between Throughput and Accuracy
Consistent with previous FLOPs-based results, Déjà Vu achieves the highest throughput-accuracy tradeoff. For instance, Déjà Vu reaches speedups of up to 1.81$\times$ for video retrieval, 2.64$\times$ for video question answering, and 2.54$\times$ for video question grounding. By contrast, Eventful Transformer, the next-best baseline, achieves speedups of 1.32$\times$, 2.08$\times$, and 2.20$\times$, respectively.
@@ 7.4. Breakdown Analysis
Figure 11. Breakdown of FLOPs on video retrieval.
@@ 7.6. Ablation Study for Inference Speedup
Figure 13. Ablation analysis compaction techniques.
We conduct an ablation study of inference speedup with a ReuseViT configuration achieving a 61% reuse rate on the video question answering task to show how each technique contributes to Déjà Vu ’s performance. Figure 13 highlights the throughput gains from each step. The first bar indicates reusing computation with hard gating alone yields a 1.25$\times$ speedup. Layer-wise sparse computation compaction then increases the speedup to 1.45$\times$ by shaping operands to better utilize the GPU kernel. Finally, layer-wise memory compaction enables larger batches, improving parallelization and raising the total speedup to 1.62$\times$. These results confirm Déjà Vu’s optimizations translate FLOPs savings into real throughput gains. Each component offers complementary benefits, gating reduces compute, sparse-compaction improves GPU workload efficiency, and memory compaction allows larger batches, unlocking significantly more performance together than individually.
7.7. Adaptability to Video Content
Figure 14. Fluctuation of computation reuse rate and cosine similarity over time.
@@ 8. Limitations and Future Work
Attention Layer Reuse. Déjà Vu currently targets only FFN and QKV layers. These layers dominate computational costs at standard resolutions, for instance, 257 tokens for ViT-L/14. However, as resolution and token counts rise—such as 577 tokens (CLIP ViT-L/14@336px), 785 tokens (DINO ViT-B/8) (Caron et al., 2021), and 1370 tokens (DINOv2 ViT-G/14@518px) (Oquab et al., 2024)—attention layers form an increasingly substantial fraction of total FLOPs, reaching up to 23.5%. Addressing this will necessitate exploring attention reuse strategies or leveraging sparse attention kernels optimized for GPUs, like those proposed in Eventful Transformer. Another approach involves specialized hardware accelerators, such as those introduced in CMC.
Broader task generalization. This work evaluates Déjà Vu specifically on retrieval, question-answering, and grounding tasks. Adapting our learned reuse strategies to other video tasks, such as action recognition or object detection, would involve adjustments to task-specific model architectures and setups. While this adaptation is conceptually straightforward, its practical exploration remains an open and promising direction for future research.
@@ 10. Conclusion
This paper addresses the computational challenges posed by ViT in the context of large-scale video analytics exploiting VideoLM. We introduce Déjà Vu, a novel algorithm-system co-designed solution, recognizing the untapped potential of exploiting a learning approach to enable inter-frame computation reuse. Central to Déjà Vu is the ReuseViT model, innovatively adapted for VideoLM tasks. This model is specifically designed for enabling computation reuse opportunities in ViT inferencing by judiciously determining the reuse targets and restoring the reuse-introducing errors automatically. We also build a ReuseViT-based Déjà Vu system, which effectively utilizes GPUs for high throughput ViT inferencing for VideoLM queries. Our evaluation shows that Déjà Vu achieves a substantial performance boost while marginally compromising the accuracy of VideoLM tasks. These findings suggest that this work represents a practical initial step toward actualizing the full potential of VideoLMs on video-language query systems.

### v2

@@ Abstract.
[…]
@@
PVLDB, 18(10): 3284 - 3298, 2025.
doi:10.14778/3748191.3748195 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 10 ISSN 2150-8097.
doi:10.14778/3748191.3748195
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/casys-kaist/DejaVu.
@@ 1. Introduction
Video data is pervasive in today’s world, playing a critical role in various applications such as autonomous machines (Deb et al., 2018; Gawande et al., 2020; Hadidi et al., 2018), education (Gan et al., 2023), healthcare (Sanderson and Matuszewski, 2022), security and surveillance (Li et al., 2019; Toor et al., 2019), e-commerce (Zhang et al., 2020a, 2022), and many others (Mun et al., 2017; Li et al., 2022d). Recently, Video-Language Models (VideoLMs) have demonstrated remarkable capabilities, offering significant potential for flexible and powerful video analytics tasks (Colas et al., 2020; Gupta and Demner-Fushman, 2022; Zhao et al., 2023; Ju et al., 2022; Yang et al., 2021, 2022a). While these powerful algorithmic advances promise enormous potential, the massive computational demand poses a significant hurdle for the widespread adoption of VideoLMs for video query systems.
@@
Figure 1 provides an overview of video-language query systems, where ViTs process video frames to extract visual embeddings, which are then fed into query-specific operations such as video retrieval, video question answering, and video question grounding. The critical role of ViTs in this process emphasizes the importance of improving ViT inference efficiency to unleash the potential of VideoLMs for building video-language query systems.
@@
ReuseViT: Reuse-enabling ViT model. We develop a modified ViT model, referred to as ReuseViT, that reuses precomputed values for similar tokens between frames. To maximize reuse opportunities, we devise a frame referencing strategy that determines inter-frame reference chains for computation reuse. ReuseViT contains Decision Layers that dynamically determine when to reuse computations by processing multiple inputs such as cosine similarity, attention scores, codec metadata, and reference types, allowing it to learn the importance and interactions of these hints automatically. Additionally, it employs Restoration Layers that learn to calibrate the reused computation values based on the changes that occurred in the current frame.
@@
[…]
@@ 3.1. Frame Selection for Computation Reuse
Reordered frame computation. To harness both past and future frames, ReuseViT processes frames in an out-of-order fashion, as shown in Figure 4(b). We categorize frames into four types, I-frames (computed independently), P-frames (referencing a previous frame), Bdist2-frames (referencing frames two steps away), and Bdist1-frames (referencing immediate neighbors), reflecting terminology akin to video codecs. Following a pattern of I $\rightarrow(P\rightarrow B\textsubscript{dist2}\rightarrow B\textsubscript{dist1}\rightarrow B\textsubscript{dist1})\rightarrow$… helps capture bidirectional temporal redundancies that purely sequential schemes may overlook. By referencing both directions, we increase the potential for reuse and reduce the frequency of full computations. When choosing references, ReuseViT evaluates the quality and temporal distance of candidate frames to preserve accuracy without incurring excessive overhead. Subsequent sections detail this decision-making process.
@@ 3.2. Layer Selection for Reuse
Figure 5. FLOPs breakdown of core computations within a single encoder layer of vision transformers at different scales.
@@
Computation patterns of layers. Besides computational cost, we consider computation patterns of each layer. The QKV projection and FFN operations are applied independently to each token, without inter-token dependencies, suiting them for computation reuse. In contrast, the self-attention operation involves interactions among all tokens, making reuse more challenging and potentially affecting model output. Given these considerations, we focus on reusing computations in the QKV projection and FFN layers, which are the most computationally intensive and token-independent.
@@ 3.3. Criteria for Computations Reuse
(4) $\mathcal{M}_{i}$ $=\begin{cases}1,&\text{if }d_{i}>0\\ 0,&\text{otherwise}\end{cases},\quad i=1,\ldots,N$
@@
[…]
@@ 6.3. Online Inference
Once deployed, Déjà Vu employs layerwise scheduling and memory compaction to translate the FLOP savings from ReuseViT into practical GPU throughput. While ReuseViT reduces the token-level computational load, effective scheduling and compaction routines ensure these savings materialize in actual GPU execution time, which is a key requirement for large-scale query systems.
@@ 7.1. Methodology
End tasks. To evaluate Déjà Vu’s effectiveness, we select three distinct Video-Language Models (VideoLMs), each for a different task. All models use CLIP (Radford et al., 2021) (ViT) embeddings as their visual backbone. We focus on:
  • Video retrieval. We use CLIP4Clip (Luo et al., 2022) on the MSR-VTT (Xu et al., 2016) dataset, which contains 10,000 text-annotated video clips of 10 to 30 seconds in length, and report top-5 recall.
  • Video question answering. For video QA, we use FrozenBiLM (Yang et al., 2022a) on How2QA (Li et al., 2020a), which has 44,007 QA pairs on 60-second clips, and report multiple-choice accuracy.
  • Video Question Grounding. For video question grounding, which requires localizing a temporal segment for an answer, we use TempCLIP (Xiao et al., 2024) on NExT-GQA (Xiao et al., 2024), with videos averaging 45 seconds. We report GQA@Accuracy, measuring correct answers with correct temporal overlap.
For a fair comparison, all methods sample frames at 2 FPS, aligning with common practice (Piergiovanni et al., 2023; Arnab et al., 2021; Bulat et al., 2021; Zhang et al., 2023; Wang et al., 2024a; Luo et al., 2022; Yang et al., 2022a; Xiao et al., 2024). We use standard dataset splits for training and evaluation for all tasks.
Baselines. As no prior methods directly accelerate VideoLMs, our baselines comprise conceptually similar temporal ViT acceleration approaches and a representative image-based method for contrast. We evaluate three acceleration methods:
  • CMC. CMC (Song et al., 2024) identifies tokens for reuse via a fixed mean squared error (MSE) threshold. Originally designed for action recognition, it can leverage specialized video codec hardware for its similarity computations.
  • Eventful Transformer. Eventful Transformer (Dutson et al., 2023) employs a simple but rigid static policy, recomputing a fixed number of tokens per layer regardless of video content. While simple and efficient, this non-adaptive strategy can cause error accumulation in dynamic scenes.
  • DiffRate. DiffRate (Chen et al., 2023) is an image-based baseline, adapted for VLP models, that combines token pruning and merging. It automatically determines the accuracy-efficiency tradeoff through backpropagation. Though originally evaluated on image classification, we adapted the policy for VLP models.
@@ 7.3. Tradeoff between Throughput and Accuracy
[…]
@@ 7.4. Breakdown Analysis
Figure 11. Breakdown of FLOPs on video retrieval. (a) FLOPs breakdown at fixed reuse rate. (b) Reuse rates comparison at equal accuracy.
@@ 7.6. Ablation Study for Inference Speedup
Figure 13. Ablation study of Déjà Vu’s inference optimization techniques on the video QA task.
[…]
@@ 7.7. Ablation Study for Design Choices
7.8. Adaptability to Video Content
Figure 15. Comparison of Déjà Vu’s adaptive reuse strategy versus Eventful Transformer’s static strategy on a video segment from How2QA over time.
@@ 8. Limitations and Future Work
Attention Layer Reuse. Déjà Vu currently targets only FFN and QKV layers. These layers dominate computational costs at standard resolutions, for instance, 257 tokens for ViT-L/14. However, as resolution and token counts rise, such as 577 tokens (CLIP ViT-L/14 at 336px), 785 tokens (DINO ViT-B/8 (Caron et al., 2021)), and 1370 tokens (DINOv2 ViT-G/14 at 518px (Oquab et al., 2024)), attention layers form an increasingly substantial fraction of total FLOPs, reaching up to 23.5%. Addressing this requires exploring attention reuse strategies or leveraging sparse attention kernels optimized for GPUs, like those proposed in Eventful Transformer. Another approach involves specialized hardware accelerators, such as those introduced in CMC.
Broader task generalization. This work evaluates Déjà Vu specifically on retrieval, question answering, and grounding tasks. Adapting our learned reuse strategies to other video tasks, such as action recognition or object detection, would involve adjustments to task-specific model architectures and setups. While this adaptation is conceptually straightforward, its practical exploration remains an open and promising direction for future research.
@@ 10. Conclusion
This paper addresses the computational challenges posed by ViT in the context of large-scale video analytics exploiting VideoLMs. We introduce Déjà Vu, a novel algorithm-system co-designed solution, recognizing the untapped potential of exploiting a learning approach to enable inter-frame computation reuse. Central to Déjà Vu is the ReuseViT model, innovatively adapted for VideoLM tasks. This model is specifically designed for enabling computation reuse opportunities in ViT inferencing by judiciously determining the reuse targets and restoring the reuse-introduced errors automatically. We also build a ReuseViT-based Déjà Vu system, which effectively utilizes GPUs for high throughput ViT inferencing for VideoLM queries. Our evaluation shows that Déjà Vu achieves a substantial performance boost while marginally compromising the accuracy of VideoLM tasks. These findings suggest that this work represents a practical initial step toward actualizing the full potential of VideoLMs for video-language query systems.
Acknowledgements.
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT), through the Information Technology Research Center (ITRC) program (No. IITP-2025-RS-2020-II201795), the Development of ML compiler framework for on-device AI (No. RS-2024-00459797), and the Graduate School of Artificial Intelligence Semiconductor program (No. IITP-2025-RS-2023-00256472).